OpenStudy (anonymous):

Differentiate (x^2-4x+2)(5x^3-x^2+5)

OpenStudy (anonymous):

\[Differentiate (x ^{2}-4x+2)(5x ^{3}-x ^{2}+5)\]

OpenStudy (anonymous):

product rule for this one use \[(fg)'=f'g+g'f\] with \[f(x)=x^2-4x+2, f'(x)=2x-4,g(x)=5x^3-x^2+5,g'(x)=15x^2-2x\]

OpenStudy (anonymous):

It says the answer should be \[25x ^{4}-84x ^{3}+42x ^{2}+6x-20\]

OpenStudy (anonymous):

f=x2-4x+2 g=5x3-x2+5 f'=2x-4 g'=15x2-2 multiply gf'+fg' answer will be 25x4-84x3+42x2+6x-20

OpenStudy (anonymous):

How did you get f=2x+4 and g=15x2-2 ?

OpenStudy (anonymous):

Can you break it down so I can try to understand? I'm sorry, I'm having a hard time. :(

OpenStudy (anonymous):

if u differentiate ax^y then you get ayx^y-1

OpenStudy (anonymous):

so x^2 when differentiated u get 2x

OpenStudy (anonymous):

How do you differentiate it from \[x ^{2}-4x+2\] to that 2x+4 ? How do you diffferentiate it from \[5x ^{3}-x ^{2}+5\] to \[15x ^{2}-2\] ?